$SPAD/src/input as-eg3.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)compile hilbert.as
monomial l == (l::Vector SingleInteger) pretend Monomial
mon1 := monomial [4,0,0,0];
mon2:= monomial [3,3,0,0];
mon3 := monomial [3,2,1,0];
mon4 := monomial[3,1,2,0];
mon5 := monomial[0,2,0,1];
mon6 := monomial[0,1,0,5];
l := [mon1, mon2, mon3, mon4, mon5, mon6];
Hilbert l
idA := varMonomsPower(6,5);
#idA
Hilbert idA
idB := varMonomsPower(6,6);
#idB
Hilbert idB
idC := varMonomsPower(12,3);
#idC
Hilbert idC
idD:=[monomial[2,0,0,0],monomial[1,1,0,0],monomial[1,0,1,0],monomial[1,0,0,1],_
 monomial[0,3,0,0],monomial[0,2,1,0]]^4;
#idD
Hilbert idD

———-

References^{[1] nothing}

4